OpenStudy (anonymous):

How can I find the derivative of tan(5/x) ?

OpenStudy (irishboy123):

if you know the derivative of tan u and the derivative of 5/x, you can use the chain rule.

OpenStudy (irishboy123):

by letting u = 5/x

OpenStudy (anonymous):

I don't know the derivative of tan u xD Is there any way to find it without having it memorized?

OpenStudy (solomonzelman):

no need for u

OpenStudy (irishboy123):

it's a really good one to know, especially for integration

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{d}{dx}\left(~\tan(x/5)~\right) }\)

OpenStudy (solomonzelman):

how would you differentiate if it was just tan(x) ?

OpenStudy (anonymous):

I don't know lol I get confused when it comes to tan

OpenStudy (solomonzelman):

what is the derivative of tan(x), you don't know ?

OpenStudy (anonymous):

Nope xD

OpenStudy (solomonzelman):

you can use a quotient rule, do you know what a quotient rule is?

OpenStudy (irishboy123):

@Tracy96 cheers

OpenStudy (solomonzelman):

Do you know what d/dx means ?

OpenStudy (anonymous):

Ohhh now I remember it was sec^2x hahaha

OpenStudy (anonymous):

Working on math too much today XD lol

OpenStudy (solomonzelman):

yes, derivative of tan(x) is sec^2(x)

OpenStudy (solomonzelman):

And when you have tan(x/5) the only difference is that it is sec^2( x/5 ), AND you need to multiply that times the chain rule..... (chain rule for x/5)

OpenStudy (solomonzelman):

For example, \(\large\color{black}{ \displaystyle \frac{d}{dx}\left(~\tan (4x)~\right)=\sec^2(4x)\times \left(\frac{d}{dx}~4x\right) }\) \(\large\color{black}{ \displaystyle \frac{d}{dx}\left(~\tan (4x)~\right)=\sec^2(4x)\times ~4 }\) \(\large\color{black}{ \displaystyle \frac{d}{dx}\left(~\tan (4x)~\right)=4\sec^2(4x) }\)

OpenStudy (anonymous):

Ok so that will be \[\sec^2(5/x)(\frac{ -5 }{ x^2 })\]

OpenStudy (solomonzelman):

yes, that is correct, but I would put the last part in front to avoid any confusion

OpenStudy (solomonzelman):

confusion*

OpenStudy (anonymous):

Right! Thanks so much @SolomonZelman :D

OpenStudy (solomonzelman):

Sure. Any questions you have regarding this problem or any of the rules?

OpenStudy (anonymous):

Nope! Thanks! That was very helpful :)

OpenStudy (solomonzelman):

Alright. You are always welcome!

OpenStudy (anonymous):

:)